OpenStudy (anonymous):

-20 degrees into radians

OpenStudy (timaashorty):

http://www.beg.utexas.edu/GIS/tools/deg_rad.htm

OpenStudy (anonymous):

-0.34906585 radians

OpenStudy (timaashorty):

enter the degrees and click calculate [:

OpenStudy (timaashorty):

-20 degrees = -0.34906585 radians

OpenStudy (anonymous):

\[180^{\circ}=\pi~radians.\] \[1^{\circ}=\frac{ \pi }{180 }radians\ \]] \[-20^{\circ}=\frac{ \pi }{180}\times \left( -20 \right)=?~radians.\]